OpenStudy (anonymous):

name the property of given statement? 2+4=2+4

OpenStudy (anonymous):

Commutative Property for Addition states that a + b = b+ a Identity Property states that a + b = a + b.

OpenStudy (anonymous):

or is it addition property of equality?

OpenStudy (anonymous):

Addition Property of Equality states that "If a = b, then a+c = b + c".

OpenStudy (anonymous):

so that's the answer?

OpenStudy (anonymous):

Identity Property.

OpenStudy (anonymous):

thats not a choice

OpenStudy (calculusxy):

Can you give the options?

OpenStudy (anonymous):

What are the choices @lisat

OpenStudy (anonymous):

What are the choices? Idenity Property is used in many texts.

OpenStudy (anonymous):

I gave you a medal @Easyaspi314

OpenStudy (anonymous):

Thanks, but why, Jazzy?

OpenStudy (anonymous):

a) addition property of equality b)distributive property c)multiplicative property of equality d)subtraction property of equality

OpenStudy (calculusxy):

a

OpenStudy (anonymous):

None of the above.

OpenStudy (anonymous):

lol okay..

OpenStudy (anonymous):

You did have the same answer as @calculusxy but you explain it to @lisat @Easyaspi314

OpenStudy (anonymous):

Well, I'm not going to change my answer. Addition Property of equality states that you can add equals to both sides of an equation. Thats NOT what you are doing. You have a+ b + a + b...that is verything is equal to itself...Identity Property.

OpenStudy (anonymous):

Is this your homework or a quiz? @lisat

OpenStudy (anonymous):

Thanks, Jazzy.

OpenStudy (anonymous):

You are welcome @Easyaspi314

OpenStudy (anonymous):

quiz

OpenStudy (anonymous):

In Geometry proofs, when a statement sich as "AB = AB", the reason thats given is Identity (or anything that means identity). This is the exact same statement just as AB = AB.

OpenStudy (anonymous):

At least your honest about it but its okay

OpenStudy (anonymous):

alright im just going with a

OpenStudy (anonymous):

I would too

OpenStudy (anonymous):

Can't be c or d because your not multiply or subtracting

OpenStudy (anonymous):

thanks guys!

OpenStudy (anonymous):

I forgot what was distributive property and You are Welcome @lisat

OpenStudy (anonymous):

@JazzyMLee9 Distributive Property: a(b + c) = ab + ac

OpenStudy (anonymous):

When 5 is added to 2 times a number, the result is 45. Find the number.

OpenStudy (anonymous):

Oh now I remember @Easyaspi314